OpenStudy (haileyd):

Can someone help and or explain how to find the vertex of the parabola by completing the square? y=x^2+8x-18

zepdrix (zepdrix):

\[\Large\rm y=\color{orangered}{x^2+8x}\qquad\qquad-18\]Hey Haily c: Ignore the -18 for now, we don't care about that guy. We want to complete the square on these x's. More generally, we think of it like this:\[\Large\rm y=\color{orangered}{x^2+bx}\]We take half of our b term, and square it. \[\Large\rm \left(\frac{b}{2}\right)^2\]That's what will complete the square for us. We'll add AND subtract that value from the right side.\[\Large\rm y=\color{orangered}{x^2+bx+\left(\frac{b}{2}\right)^2}-\left(\frac{b}{2}\right)^2\]The orange part condenses down to a perfect square,\[\Large\rm y=\color{orangered}{\left(x+\frac{b}{2}\right)^2}-\left(\frac{b}{2}\right)^2\]

zepdrix (zepdrix):

Maybe a confusing explanation ^ Let's just jump into it and see how it goes. So we'll take half of the b coefficient, half of 8 is 4, yah? And square it. \(\Large\rm \left(\dfrac{8}{2}\right)^2=16\) This is the value that will complete the square for us.

zepdrix (zepdrix):

\[\Large\rm y=\color{orangered}{x^2+8x+16}\qquad\qquad-18\]So we want to add 16 to complete our square. But we can't just throw in a 16 willy nilly like that. We have to balance it by subtracting 16 at the same time.\[\Large\rm y=\color{orangered}{x^2+8x+16}-16-18\]

zepdrix (zepdrix):

Then the orange part will condense down, that is our perfect square. What do you think Hailey? :d you with me?

OpenStudy (dtan5457):

Also, if you just want the vertex and you weren't forced to complete the square. You would just do -b/2a That would the x coordinate of your vertex. Plug that back into x^2+8x-18 for the y value. otherwise follow @zepdrix

OpenStudy (haileyd):

Im still reading over the explanations so just a few more seconds and ill be with you.

OpenStudy (haileyd):

Okay i think i got it for here thanks:)

zepdrix (zepdrix):

Got it from there? c: k